X, Y, Z are three sets in a sample space S. Find P(X|Y ∩Z) if P(X|Y ) = 0.1
and P(X|Z) = 0.35 are given.
Answer #1

We are given here that:
P(X | Y) = 0.1,
P(X | Z) = 0.35

Given that both Y and Z has happened, and assuming that Y and Z are independent events here, we use the multiplication rule to get the probability here as:
P(X|Y ∩Z) = P(X | Y)P(X | Z) = 0.1*0.35 = 0.035

Therefore 0.035 is the required probability here.